# AlphaDecay: Module-wise Weight Decay for Heavy-Tailed Balancing in LLMs

### v2

@@ Abstract
Weight decay is a standard regularization technique for training large language models (LLMs). While it is common to assign a uniform decay rate to every layer, this approach overlooks the structural diversity of LLMs and the varying spectral properties across modules. In this paper, we introduce AlphaDecay, a simple yet effective method that adaptively assigns different weight decay strengths to each module of an LLM. Our approach is guided by Heavy-Tailed Self-Regularization (HT-SR) theory, which analyzes the empirical spectral density (ESD) of weight correlation matrices to quantify “heavy-tailedness.” Modules exhibiting more pronounced heavy-tailed ESDs, reflecting stronger feature learning, are assigned weaker decay, while modules with lighter-tailed spectra receive stronger decay. Our method leverages tailored weight decay assignments to balance the module-wise differences in spectral properties, leading to improved performance. Extensive pre-training tasks with various model sizes from 60M to 1B demonstrate that AlphaDecay achieves better perplexity and generalization than conventional uniform decay and other adaptive decay baselines. Our code is available at https://github.com/hed-ucas/AlphaDecay.
@@ 1 Introduction
Large language models (LLMs) have emerged as a core technology in artificial intelligence, with extensive applications in chatbots, content generation, code synthesis, and other domains, significantly enhancing the efficiency and user experience of human-computer interaction [5; 29; 23; 11]. However, the formidable capabilities of these models rely on massive pre-training datasets and substantial computational resources, rendering the training process fraught with challenges [45; 2]. Persistent research challenges include, but are not limited to, the efficient optimization of ultra-large-scale parameters and the trade-off between training costs and model performance.
@@
Weight decay, one of the most widely used regularization techniques for training well-generalized deep neural networks [26; 43; 13], critically influences the convergence and performance of state-of-the-art machine learning algorithms when properly configured. Extensive prior studies [21; 33; 36] have demonstrated its pivotal role in enhancing model generalization from diverse theoretical and empirical perspectives. Recent work [20; 8] further highlights its importance in improving optimizer stability and efficacy during the training of LLMs.
@@
The prevailing approach to weight decay assigns a globally fixed value per epoch across optimizers—including SGD [40], Adam [18], and their variants [48; 37]—where all model layers share an identical decay coefficient. However, given the scaling parameter counts and architectural complexity of modern LLMs, such a uniform weight decay scheme fails to capture their intricate structural properties, making this conventional practice increasingly suboptimal. Notably, recent work has begun investigating dynamic weight decay adaptation [16; 31; 12; 43] to address this limitation. [12] observes that fixed-hyperparameter weight decay fails to balance robustness and accuracy in adversarial training, causing robust overfitting. They propose Adaptive Weight Decay (AWD) to dynamically adjust decay strength via classification and regularization loss gradients, automatically enhancing robustness and adversarial performance without extra data.
@@
Figure 1: Module-wise Balance and AlphaDecay weight decay schedule. (a) Employing PL fitting to derive module-wise PL_Alpha_Hill values, AlphaDecay achieves module-wise balance by increasing the lower values (e.g., att.Q and att.K, more heavy-tailed) while decreasing the higher values (e.g., MLP components, less heavy-tailed). (b) Given the imbalanced module-wise PL_Alpha_Hill of LLaMa-60M, AlphaDecay assigns lower weight decay to modules with lower PL_Alpha_Hill.
@@
Notably, prior studies on dynamic weight decay adaptation were exclusively designed for architectures like ResNet18/34/50 [14], VGG [35], and DenseNet [15], employing time-wise modulation (i.e., uniform decay values across all layers at each timestep) while maintaining layer-wise uniformity. This approach is reasonable for parameter-efficient, structurally simple models (e.g., ResNets) where inter-layer feature distinctions are less pronounced. However,
@@
Three reasons behoove us to pose the above research question: First, the prevailing consensus holds that certain transformer components exhibit greater functional importance than others [41; 4; 45; 27], necessitating differentiated weight decay treatment. Second, weight decay manifests fundamentally distinct roles in over-trained regimes (e.g., ResNets) versus under-trained regimes (e.g., LLMs) [8]. Most notably, existing research demonstrates that improper weight decay configuration for LLMs may adversely affect model performance [3; 19; 43; 34; 17; 9]. Our main contributions are as follows:
@@
  • ❸ Extensive experiments spanning models from 60M to 1B parameters show that the proposed approach, AlphaDecay, consistently outperforms the Uniform baseline as well as adaptive methods such as AWD [12] and AdaDecay [31] (see table 2). These results highlight the critical role of module-wise balance in achieving state-of-the-art performance in LLMs.
@@ 2 Related Work
Weight decay in LLM training. Weight decay is a widely adopted technique for training deep networks, spanning applications from image classification to LLMs [21]. In the context of GPT-3 training, [5] recommended incorporating weight decay primarily for its mild regularization benefits. [20] showed that weight decay promotes optimizer equilibrium in scale-invariant systems. Recent studies have provided deeper insights into weight decay’s role in LLM training. [39; 8; 38] challenged the conventional view of weight decay’s generalization benefits for LLMs, and instead highlighting its critical function in reducing training loss and enhancing stability during under-training through the lens of Effective Learning Rate. Building on these findings, [3; 19] established a connection between $l_{2}$ regularization and spectral norms, discovering that weight decay induces low-rank attention layers. Their work further showed that employing different weight decay values for attention and MLP modules, carefully tuned via grid search, can significantly improve training outcomes. Our work presents the first formal analysis of non-uniform module-wise weight decay in LLM training, demonstrating its effectiveness through comprehensive empirical validation.
@@
Dynamic weight decay. While uniform weight decay is commonly used for model training, a line of work employs gradnorm to adaptively determine weight decay settings. [16] analyzed gradient descent with weight decay, finding that backpropagated gradients scale with upstream weights while weight decay scales with each layer’s own weights. This mismatch in scaling causes layer-wise overfitting or underfitting, leading them to propose using the gradient-to-decay magnitude ratio as a layer-wise coefficient. [31] enhanced this approach by normalizing gradients and applying a scaled sigmoid to compute the coefficient. Similarly, [12] used the ratio of gradient norms to parameter norms. [43] showed weight decay amplifies late-stage gradient norms, harming convergence. Their solution, AdamS, penalizes large gradients and outperforms both Adam and AdamW. Another line of research [3; 10; 42] revealed that weight decay induces low-rank layer structures. [19] further showed that applying distinct weight decay values to attention and MLP modules, meticulously tuned via grid search, can substantially enhance training outcomes. Building upon these foundations, our work advances this direction by introducing the first weight decay scheduling framework for LLMs.
@@
Heavy-tail self-regularization. HR-SR Theory examines the ESD of weight matrices and identifies its relationship with training quality based on principles from statistical physics and random matrix theory [7]. HT-SR Theory posits that well-trained neural networks exhibit strong correlations in their weights, manifesting as heavy-tailed structures in the ESD of each layer’s weight matrics [28; 30]. Recently, HT-SR has been applied to model selection [28; 30; 44], module-wise adaptive training [47], and LLM pruning [27], demonstrating its efficacy in estimating model and layer quality. However, no prior work has explored HR-SR theory in the context of weight decay configuration. Our work draws inspiration from HR-SR theory and introduces a novel technique that leverages ESD structures to guide weight decay settings.
@@ 3.1 Rationale
Different modules in LLMs exhibit diverse spectral properties, particularly in the distribution of their singular values. Figure 2 visualize the normalized singular value spectra of the weight matrices for each module type (att.q, att.k, att.v, att.o, mlp.gate, mlp.up, mlp.down) across all 40 transformer layers in the pretrained LLaMa-2-13b-hf model. Notably, substantial variability is observed in the heavy-tailedness of the singular value distributions: the attention-related modules (att.q and att.k) consistently show heavier tails, while the MLP modules (mlp.gate, mlp.up, mlp.down) exhibit lighter tails (see figure 3).
@@
Figure 3: Comparison of ESD distributions across modules of LLaMa-135M under different training methods (AlphaDecay vs. Uniform). Attention-related modules (e.g., att.q, att.k) exhibit notably heavier spectral tails in contrast to MLP-associated modules. Our method systematically balances the heavy-tailed properties across modules by appropriately configuring module-wise weight decay, thereby enhancing overall model performance.
3.2 HR-SR Theory
@@
Given a network with $N$ modules and weight matrices $\{\mathbf{W_{l}}\}_{l=1}^{L}$ of shape $n\times m$ ($n\leq m$), we compute the ESD by obtaining the eigenvalues of the correlation matrix $\mathbf{X}_{l}=\mathbf{W}_{l}^{\top}\mathbf{W}_{l}$ for each module. The power law fit for the ESD takes the form:
@@
To estimate $\alpha$, we use the Hill estimator [47; 25]. For a given module’s eigenvalues $\{\lambda_{i}\}_{i=1}^{n}$ (sorted in ascending order), the Hill estimator is given by:
[…]
@@
PL_Alpha_Hill is a key spectral descriptor for analyzing model performance. Related works [47; 25] suggest that lower PL_Alpha_Hill values indicate "overtrained" layers (compared to other layers in the model), while higher values indicate "undertrained" layers. An important conclusion is that a more uniform distribution of PL_Alpha_Hill across layers reflects more balanced training quality, leading to better overall model quality. While these findings highlight the importance of layer-wise training balance, our work emphasizes a complementary perspective:
@@ 3.3 AlphaDecay
where $\eta$ is the initial weight decay, and $(s_{1},s_{2})$ define the range of scaling ratios applied to $\eta$. $\alpha_{t}^{i}$ is the PL_Alpha_Hill value of module $i$ at step $t$, while $\alpha_{t}^{min}$ and $\alpha_{t}^{max}$ are the minimum and maximum PL_Alpha_Hill values among all modules at step $t$. Formula (3) guarantees that the adjusted weight decay, $f_{t}(i)$, remains within $[s_{1}\eta,s_{2}\eta]$ as a scaled variant of $\eta$.
@@
             Compute $\alpha_{t}^{i}$ for all modules using the Hill estimator (2);
             Leverage all $\alpha_{t}^{i}$ and adopt $f_{t}$ in (3) to assign module-wise weight decay between $s_{1}\eta$ and $s_{2}\eta$;
@@ 4.1 Experimental setup
Models and Datasets. We conduct a systematic evaluation of AlphaDecay across LLaMa-based architectures spanning four model scales (60M, 135M, 350M, and 1B parameters). All experiments employ the C4 dataset [32], a rigorously processed subset of Common Crawl widely adopted for language model pretraining. Our experimental design incorporates two key components: (1) a non-repeating data regime with sufficient tokens for convergence, and (2) standardized preprocessing pipelines across all model scales. This multi-scale approach facilitates systematic comparison of model behaviors across different capacity regimes, while minimizing potential confounding factors in the analysis.
@@ 4.2 LLM Pre-training
LLaMa-60M LLaMa-135M LLaMa-350M LLaMa-1B
Weight Decay 1e-5 5e-6 1e-6 1e-5 5e-6 1e-6 1e-5 5e-6 1e-6 1e-5 5e-6 1e-6
WD=0 33.23 24.60 18.62 16.11
Uniform 32.39 32.56 33.03 22.99 23.14 24.14 17.12 16.74 17.50 15.36 14.66 15.03
AWD[12] 33.78 33.74 33.74 24.25 24.45 24.53 18.32 18.55 18.79 16.03 16.22 16.38
Adadecay[31] 32.24 32.52 33.03 23.20 23.08 23.96 18.21 17.42 17.91 17.23 18.14 15.35
AlphaDecay 31.56 31.58 32.61 22.76 22.55 23.49 17.00 16.66 16.88 15.13 14.55 14.63
Table 2 presents the main results of our study, where we evaluate the effectiveness of different weight decay scheduling strategies on the pre-training of LLaMa models with varying parameter scales (60M, 135M, 350M, and 1B) on the C4 dataset. For each model size, we conduct comprehensive experiments across three commonly used weight decay values (1e-5, 5e-6, and 1e-6). Our proposed method is compared against several baselines, including the commonly used Uniform scheduling, adaptive global weight decay (AWD) [12], and adaptive per-module weight decay (Adadecay) [31]. All baseline methods are carefully tuned for a fair comparison.

### v3

@@ Abstract
Weight decay is a standard regularization technique for training large language models (LLMs). While it is common to assign a uniform decay rate to every layer, this approach overlooks the structural diversity of LLMs and the varying spectral properties across modules. In this paper, we introduce AlphaDecay, a simple yet effective method that adaptively assigns different weight decay strengths to each module of an LLM. Our approach is guided by Heavy-Tailed Self-Regularization (HT-SR) theory, which analyzes the empirical spectral density (ESD) of weight correlation matrices to quantify “heavy-tailedness.” Modules exhibiting more pronounced heavy-tailed ESDs, reflecting stronger feature learning, are assigned weaker decay, while modules with lighter-tailed spectra receive stronger decay. Our method leverages tailored weight decay assignments to balance the module-wise differences in spectral properties, leading to improved performance. Extensive pre-training tasks with various model sizes from 60M to 1B demonstrate that AlphaDecay achieves better perplexity and generalization than conventional uniform decay and other adaptive decay baselines. The code is available at https://github.com/hed-ucas/AlphaDecay.
@@ 1 Introduction
Large language models (LLMs) have emerged as a core technology in artificial intelligence, with extensive applications in chatbots, content generation, code synthesis, and other domains, significantly enhancing the efficiency and user experience of human-computer interaction [brown2020language; maini2024llm; liu2021we; ge2023openagi]. However, the formidable capabilities of these models rely on massive pre-training datasets and substantial computational resources, rendering the training process fraught with challenges [yin2023outlier; an2024make]. Persistent research challenges include, but are not limited to, the efficient optimization of ultra-large-scale parameters and the trade-off between training costs and model performance.
@@
Weight decay, one of the most widely used regularization techniques for training well-generalized deep neural networks [loshchilov2017decoupled; xie2023overlooked; hardt2016train], critically influences the convergence and performance of state-of-the-art machine learning algorithms when properly configured. Extensive prior studies [krogh1991simple; shalev2014understanding; soudry2018implicit] have demonstrated its pivotal role in enhancing model generalization from diverse theoretical and empirical perspectives. Recent work [kosson2023rotational; d2024we] further highlights its importance in improving optimizer stability and efficacy during the training of LLMs.
@@
The prevailing approach to weight decay assigns a globally fixed value per epoch across optimizers—including SGD [wang2021cooperative], Adam [kingma2014adam], and their variants [zhuang2020adabelief; stich2018sparsified]—where all model layers share an identical decay coefficient. However, given the scaling parameter counts and architectural complexity of modern LLMs, such a uniform weight decay scheme fails to capture their intricate structural properties, making this conventional practice increasingly suboptimal. Notably, recent work has begun investigating dynamic weight decay adaptation [ishii2017layer; nakamura2019adaptive; ghiasi2023improving; xie2023overlooked] to address this limitation. [ghiasi2023improving] observes that fixed-hyperparameter weight decay fails to balance robustness and accuracy in adversarial training, causing robust overfitting. They propose Adaptive Weight Decay (AWD) to dynamically adjust decay strength via classification and regularization loss gradients, automatically enhancing robustness and adversarial performance without extra data.
@@
Figure 1: Module-wise Balance and AlphaDecay weight decay schedule. (a) Employing PL fitting to derive module-wise PL_Alpha_Hill values (see formula (2)), AlphaDecay achieves module-wise balance by increasing the lower values (e.g., att.Q and att.K, more heavy-tailed) while decreasing the higher values (e.g., MLP components, less heavy-tailed). (b) Given the imbalanced module-wise PL_Alpha_Hill of LLaMa-60M, AlphaDecay assigns lower weight decay to modules with lower PL_Alpha_Hill.
@@
Notably, prior studies on dynamic weight decay adaptation were exclusively designed for architectures like ResNet18/34/50 [he2016deep], VGG [simonyan2014very], and DenseNet [huang2017densely], employing time-wise modulation (i.e., uniform decay values across all layers at each timestep) while maintaining layer-wise uniformity. This approach is reasonable for parameter-efficient, structurally simple models (e.g., ResNets) where inter-layer feature distinctions are less pronounced. However,
@@
Three reasons behoove us to pose the above research question: First, the prevailing consensus holds that certain transformer components exhibit greater functional importance than others [wang2020rethinking; bhojanapalli2021leveraging; yin2023outlier; lu2024alphapruning], necessitating differentiated weight decay treatment. Second, weight decay manifests fundamentally distinct roles in over-trained regimes (e.g., ResNets) versus under-trained regimes (e.g., LLMs) [d2024we]. Most notably, existing research demonstrates that improper weight decay configuration for LLMs may adversely affect model performance [bhojanapalli2020low; kobayashi2024weight; xie2023overlooked; sharma2023truth; jaiswal2024galore; fernandez2025full]. Our main contributions are as follows:
@@
  • ❸Extensive experiments spanning models from 60M to 1B parameters show that the proposed approach, AlphaDecay, consistently outperforms the Uniform baseline as well as adaptive methods such as AWD [ghiasi2023improving] and AdaDecay [nakamura2019adaptive] (see table 2). These results highlight the critical role of module-wise balance in achieving state-of-the-art performance in LLMs.
@@ 2 Related Work
Weight decay in LLM training. Weight decay is a widely adopted technique for training deep networks, spanning applications from image classification to LLMs [krogh1991simple]. In the context of GPT-3 training, [brown2020language] recommended incorporating weight decay primarily for its mild regularization benefits. [kosson2023rotational] showed that weight decay promotes optimizer equilibrium in scale-invariant systems. Recent studies have provided deeper insights into weight decay’s role in LLM training. [van2017l2; d2024we; tu2024online] challenged the conventional view of weight decay’s generalization benefits for LLMs, and instead highlighting its critical function in reducing training loss and enhancing stability during under-training through the lens of Effective Learning Rate. Building on these findings, [bhojanapalli2020low; kobayashi2024weight] established a connection between $l_{2}$ regularization and spectral norms, discovering that weight decay induces low-rank attention layers. Their work further showed that employing different weight decay values for attention and MLP modules, carefully tuned via grid search, can significantly improve training outcomes. Our work presents the first formal analysis of non-uniform module-wise weight decay in LLM training, demonstrating its effectiveness through comprehensive empirical validation.
@@
Dynamic weight decay. While uniform weight decay is commonly used for model training, a line of work employs gradnorm to adaptively determine weight decay settings. [ishii2017layer] analyzed gradient descent with weight decay, finding that backpropagated gradients scale with upstream weights while weight decay scales with each layer’s own weights. This mismatch in scaling causes layer-wise overfitting or underfitting, leading them to propose using the gradient-to-decay magnitude ratio as a layer-wise coefficient. [nakamura2019adaptive] enhanced this approach by normalizing gradients and applying a scaled sigmoid to compute the coefficient. Similarly, [ghiasi2023improving] used the ratio of gradient norms to parameter norms. [xie2023overlooked] showed weight decay amplifies late-stage gradient norms, harming convergence. Their solution, AdamS, penalizes large gradients and outperforms both Adam and AdamW. Another line of research [bhojanapalli2020low; galanti2022characterizing; wang2023implicit] revealed that weight decay induces low-rank layer structures. [kobayashi2024weight] further showed that applying distinct weight decay values to attention and MLP modules, meticulously tuned via grid search, can substantially enhance training outcomes. Building upon these foundations, our work advances this direction by introducing the first weight decay scheduling framework for LLMs.
@@
Heavy-tail self-regularization. HT-SR Theory examines the ESD of weight matrices and identifies its relationship with training quality based on principles from statistical physics and random matrix theory [couillet2022random]. HT-SR Theory posits that well-trained neural networks exhibit strong correlations in their weights, manifesting as heavy-tailed structures in the ESD of each layer’s weight matrics [mahoney2019traditional; martin2021predicting]. Recently, HT-SR has been applied to model selection [mahoney2019traditional; martin2021predicting; yang2023test], module-wise adaptive training [zhou2023temperature], and LLM pruning [lu2024alphapruning], demonstrating its efficacy in estimating model and layer quality. However, no prior work has explored HT-SR theory in the context of weight decay configuration. Our work draws inspiration from HT-SR theory and introduces a novel technique that leverages ESD structures to guide weight decay settings.
@@ 3.1 Rationale
Different modules in LLMs exhibit diverse spectral properties, particularly in the distribution of their singular values. Figure 2 visualize the normalized singular value spectra of the weight matrices for each module type (att.q, att.k, att.v, att.o, mlp.gate, mlp.up, mlp.down) across all 40 transformer layers in the pretrained LLaMa-2-13b-hf model. Notably, substantial variability is observed in the heavy-tailedness of the singular value distributions: the attention-related modules (att.q and att.k) consistently show heavier tails, while the MLP modules (mlp.gate, mlp.up, mlp.down) exhibit lighter tails.
@@
Figure 3: Comparison of ESD distributions across modules of LLaMa-135M under different training methods (AlphaDecay: Perplexity=22.55 vs. Uniform: Perplexity=23.14). Attention-related modules (e.g., att.q, att.k) exhibit notably heavier spectral tails in contrast to MLP-associated modules. Our method systematically balances the heavy-tailed properties across modules by appropriately configuring module-wise weight decay, thereby enhancing overall model performance.
3.2 HT-SR Theory
@@
Given a network with $L$ modules and weight matrices $\{\mathbf{W_{l}}\}_{l=1}^{L}$ of shape $n\times m$ ($n\leq m$), we compute the ESD by obtaining the eigenvalues of the correlation matrix $\mathbf{X}_{l}=\mathbf{W}_{l}^{\top}\mathbf{W}_{l}$ for each module. The power law fit for the ESD takes the form:
@@
To estimate $\alpha$, we use the Hill estimator [zhou2023temperature; liu2024model]. For a given module’s eigenvalues $\{\lambda_{i}\}_{i=1}^{n}$ (sorted in ascending order), the Hill estimator is given by:
$\texttt{PL_Alpha_Hill}=1+\frac{k}{\sum_{i=1}^{k}ln\frac{\lambda_{n-i+1}}{\lambda_{n-k}}}$ (2)
@@
PL_Alpha_Hill is a key spectral descriptor for analyzing model performance. Related works [zhou2023temperature; liu2024model] suggest that lower PL_Alpha_Hill values indicate "overtrained" layers (compared to other layers in the model), while higher values indicate "undertrained" layers. An important conclusion is that a more uniform distribution of PL_Alpha_Hill across layers reflects more balanced training quality, leading to better overall model quality. While these findings highlight the importance of layer-wise training balance, our work emphasizes a complementary perspective:
@@ 3.3 AlphaDecay
where $\eta$ is the initial weight decay, and $(s_{1},s_{2})$ define the range of scaling ratios applied to $\eta$. $\alpha_{t}^{i}$ is the PL_Alpha_Hill value of module $i$ at step $t$, while $\alpha_{t}^{min}$ and $\alpha_{t}^{max}$ are the minimum and maximum PL_Alpha_Hill values among all modules at step $t$. Formula (3) guarantees that the adjusted weight decay, $f_{t}(i)$, remains within $s_{1}\eta,s_{2}\eta]$ as a scaled variant of $\eta$.
@@
      Compute $\alpha_{t}^{i}$ for all modules using the Hill estimator;
      Leverage all $\alpha_{t}^{i}$ and adopt $f_{t}(i)=\eta\cdot(\frac{\alpha_{t}^{i}-\alpha_{t}^{\text{min}}}{\alpha_{t}^{\text{max}}-\alpha_{t}^{\text{min}}}(s_{2}-s_{1})+s_{1})$ to assign module-wise weight decay between $s_{1}\eta$ and $s_{2}\eta$;
@@ 4.1 Experimental setup
Models and Datasets. We conduct a systematic evaluation of AlphaDecay across LLaMa-based architectures spanning four model scales (60M, 135M, 350M, and 1B parameters). All experiments employ the C4 dataset [raffel2020exploring], a rigorously processed subset of Common Crawl widely adopted for language model pretraining. Our experimental design incorporates two key components: (1) a non-repeating data regime with sufficient tokens for convergence, and (2) standardized preprocessing pipelines across all model scales. This multi-scale approach facilitates systematic comparison of model behaviors across different capacity regimes, while minimizing potential confounding factors in the analysis.
@@ 4.2 LLM Pre-training
LLaMa-135M LLaMa-350M LLaMa-1B
Weight Decay 1e-5 5e-6 1e-6 1e-5 5e-6 1e-6 1e-5 5e-6 1e-6
WD=0 24.60 18.62 16.11
Uniform 22.99 23.14 24.14 17.12 16.74 17.50 15.36 14.66 15.03
AWD[ghiasi2023improving] 24.25 24.45 24.53 18.32 18.55 18.79 16.03 16.22 16.38
Adadecay[nakamura2019adaptive] 23.20 23.08 23.96 18.21 17.42 17.91 17.23 18.14 15.35
AlphaDecay 22.76 22.55 23.49 17.00 16.66 16.88 15.13 14.55 14.63
Table 2 presents the main results of our study, where we evaluate the effectiveness of different weight decay scheduling strategies on the pre-training of LLaMa models with varying parameter scales (60M, 135M, 350M, and 1B) on the C4 dataset. For each model size, we conduct comprehensive experiments across three commonly used weight decay values (1e-5, 5e-6, and 1e-6). Our proposed method is compared against several baselines, including the commonly used Uniform scheduling, adaptive global weight decay (AWD) [ghiasi2023improving], and adaptive per-module weight decay (Adadecay) [nakamura2019adaptive]. All baseline methods are carefully tuned for a fair comparison.
@@
4.3 Downstream tasks &\& architectures
This section introduces our evaluation of downstream gains across zero-shot commonsense reasoning and fine-tuning tasks, with results on additional model architecture and image classification task (GPT-nano/C4 perplexity; ViT-tiny/ImageNet-1K Top-1).
Zero-shot Results.
We evaluate the pretrained LLaMa‑1B checkpoints from Table 2 on seven zero-shot commonsense reasoning tasks using lm-eval-harness with its default prompts. As shown in Table 3, AlphaDecay delivers the best results on 6 of 7 benchmarks, indicating that its pretraining gains transfer well to downstream reasoning tasks and support broad applicability.
Table 3: (Zero-shot results of commonsense‑reasoning tasks). Zero-shot evaluation results ($\uparrow$) on seven commonsense reasoning benchmarks using the LLaMa-1B model pretrained with different methods.
Method ARC-c ARC-e PIQA Hellaswag OBQA Winogrande BOOLQ Avg.
Uniform 20.22 46.72 67.68 32.94 18.8 49.41 54.74 41.50
AdaDecay 19.20 46.72 66.97 32.96 18.0 51.54 56.36 41.68
AWD 19.18 46.34 66.65 31.37 18.0 51.07 57.25 41.41
AlphaDecay 20.90 48.86 68.44 34.16 19.80 50.59 60.70 43.35
Finetuning Results.
We evaluate all baselines on GLUE finetuning tasks with roberta-base. As reported in Table 4, AlphaDecay attains the top result on 6 of 7 tasks. This evidences that AlphaDecay is effective not only during pretraining but also transfers well to finetuning settings.
Table 4: (Finetuning tasks). Finetuning results ($\uparrow$) on eight benchmarks from the GLUE dataset using roberta-base with different methods.
Method cola mnli mrpc qnli qqp rte sst2 stsb Avg.
Uniform 59.73 86.78 87.01 92.59 89.97 70.11 93.69 90.78 83.83
AdaDecay 60.45 87.23 88.19 92.62 89.95 73.36 93.73 90.9 84.55
AWD 60.72 87.44 89.53 92.58 90.08 72.27 93.72 90.9 84.66
AlphaDecay 62.82 87.11 89.61 92.73 90.12 73.86 93.77 90.91 85.12
Across architectures and datasets.
We evaluate all methods on GPT-nano/C4 (perplexity) and ViT-tiny/ImageNet-1K (Top-1 accuracy). As shown in Table 5, our method attains the lowest perplexity on GPT-nano/C4 and the highest Top-1 accuracy on ViT-tiny/ImageNet-1K, outperforming Uniform, AWD, and AdaDecay. These results demonstrate consistent effectiveness across different architectures and datasets, and strong generalization beyond a single setting.
Table 5: (Across architectures and datasets). Results on GPT-nano/C4 (Perplexity) and ViT-tiny/ImageNet-1K (Top-1) with different methods.
Backbone / Dataset Metric Uniform AWD AdaDecay AlphaDecay
GPT-nano / C4 PPL($\downarrow$) 27.56 27.64 27.68 27.37
[…]